OpenStudy (anonymous):

arithmetic Sequence: (2, 4, 8, 16....) number of times is 10. Need: an=a1+(n-1)d also need the sum: Sn=n/2(a1+an) Help? I'm stuck.

OpenStudy (anonymous):

Do you have a1 yet?

OpenStudy (anonymous):

a1 should be 2 since its the first term

OpenStudy (anonymous):

its "d" that I am having a problem with....

OpenStudy (anonymous):

Yep, do you have difference d yet?

OpenStudy (anonymous):

if you take 4 -2 =2.....doesn't work for the total problem

OpenStudy (anonymous):

I believe it's not arithmetic!

OpenStudy (anonymous):

its either arithmetic or geometric sequence...since this is the worksheet that I am working on

OpenStudy (anonymous):

It's obviously geometric! Can you find ratio r?

OpenStudy (anonymous):

Ok, haven't learned this part yet....but you can explain it and I'll be ahead for Monday's class!!!! LOL

OpenStudy (anonymous):

If you want...or I can wait til Monday!

OpenStudy (anonymous):

Instead of subtract, now you divide r = 4/2 = 8/4 = ...

OpenStudy (anonymous):

and the rest of the formula is the same???

OpenStudy (anonymous):

a_10 = a_1 * r^ 9

OpenStudy (anonymous):

ok, thanks for your help...I better wait til Monday for this problem to see how my professor wants us to work this type...Thanks Again!!!

OpenStudy (anonymous):

:)